8-15

The following information describes production activities of Mercer Manufacturing for the year. Actual direct materiaïs used Actual direct labor used 32,000 Ïbs. at ș5.85 per lb. 10,610 hours for a total of $218,566 63,030 Actual units produced Budgeted standards for each unit produced are 0.50 pounds of direct material at $5.80 per pound and 10 minutes of direct labor at $21.60 per hour. AQ Actual Quantity SQ Standard Quantity AP- Actual Price SP Standard Price AH Actual Hours SH-Standard Hours AR Actual Ratee SR Standard Rate(1) Compute the direct materials price and quantity variances (2) Compute the direct labor rate and efficiency variances. Indicate whether each variance is favorable or unfavorable. Complete this question by entering your answers in the tabs below. Required 1Required 2 Compute the direct materials price and quantity variances. (Do not round intermediate calculations.) Actual Cost Standard Cost(1) Compute the direct materials price and quantity variances. (2) Compute the direct labor rate and efficiency variances. Indicate whether each variance is favorable or unfavorable. Complete this question by entering your answers in the tabs below. Required 1Required 2 Compute the direct labor rate and efficiency variances. Indicate whether each variance is favorable or unfavorable. (Do not round intermediate calculations.) Actual Cost Standard Cost
Answer #1

1.)

Actual Cost

Standard Cost

Actual qty

Actual

Price

Actual cost

Actual qty

Standard

price

Standard

Cost of

Actual qty

Standard

Quantity

Std

price

Standard

cost

32000lbs

$ 5.85

$187200

32000 lbs

$ 5.80

$ 185600

31515

$5.80

$182787

Material Price Variance

1600 U

Material Quantity Variance

2813 A

Material Price Variance

1600

U

Material Quantity Variance

2813

U

Total Material Cost Variance

4413

U

2)

Actual Cost

Standard Cost

Actual

Hours

Actual

Rate

Actual

Cost

Actual

Hours

Standard

Rate

Standard

Cost of

Actual

hours

Standard

Hours

Standard

Rate

Standard

Cost

10610

$ 20.60

$218566

10610

$21.60

$229176

10505

$21.60

$226908

Labour Rate Variance

10610 F

Labour Efficiency Variance

2268U

Labour Rate Variance

$

10610 F

Labour Efficiency Variance

$

2268 U

Total Labour Cost Variance

$

8342 F

Working Notes:

Material Price Variance=(Standard Price-Actual Price)*Actual Quantity Purchased

Material Price Variance =(5.80-5.85)*32000=1600(U)

Material Quantity Variance=(Standard Quantity-Actual Quantity)*Standard Price

Material Quantity Variance=(31515-32000)*5.80=2813(U)

Direct Labour Rate Variance=(Standard Rate/hour-Actual Rate/hour)*Actual hours

                                           =(Standard rate*Actual hours-actual rate*Actual hours)

                                            = $21.60*10610 hours-$ 218566=$ 10610(F)

Direct Labour Efficiency Variance=(Standard hours-Actual hours)* Standard rate/hour

                                                              =(10505 hours-10610 hours)*$ 21.60= $ 2268(U)